Find the altitude of a triangle whose area is 8in.^2 and whose base is twice the altitude.
Answer by stanbon(75887)

Find the altitude of a triangle whose area is 8in.^2 and whose base is twice the altitude.
---------------------
Area = (1/2)(base)(altitude)
8 = (1/2)(2x)(x)
8 = x^2
x = 2sqrt2 in. (this is the altitude)
============
Cheers,
Stan H.